# Shapley Values-Powered Framework for Fair Reward Split in Content Produced by GenAI

### v1

@@ 1 Introduction
  1. 1.
    The artist earns more compensation the closer the generation aligns with their style.
  2. 2.
    Artist receives extra compensation if the generated content is similar to content they have previously created.
  1. 1.
    Transferring the style to new content not inherent to the given artist.
  2. 2.
    Combining the styles of multiple artists.
@@ 4.1 Evaluation of a single style
Thus, we form a dataset as depicted in Figure 4. We aim to assess the extent to which the images from the second and third columns influence the first one. As previously mentioned, we evaluate this influence by measuring the similarity score between the style-less generations, the original artworks, and the final result. This evaluation produces the graph shown in Figure 4(a). From this graph, it’s evident that the original artworks by the artist bear a significant resemblance to the final generation. This observation can also be visually corroborated using Figure 3. When we input the values from the graph into the Shapley Values computation function, we obtain the following values: {’Alena Aenami’:0.7422, ’Stable Diffusion’:0.2578} (see Figure 4(b)).
@@
(a) prompt: a city
(b) prompt: a city, by Alena Aenami
@@ 4.2.1 Dreambooth
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
(a) image produced by Dreambooth of Pixel Jeff
(b) image produced by Dreambooth of Tony Skeor
(c) mix of Pixel Jeff and Tony Skeor models
@@
(a) image produced by Dreambooth of Greg Rutkowski
(b) image produced by Dreambooth of Leon Tukker
(c) mix of Greg Rutkowski and Leon Tukker models
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
(a) image produced by Dreambooth of Greg Rutkowski
(b) image produced by Dreambooth of Leon Tukker
(c) mix of Greg Rutkowski and Leon Tukker models
(d) outputs from the SD1.5 model without any specific style applied.
@@
(a) image produced by Dreambooth of Greg Rutkowski
(b) image produced by Dreambooth of Even Amundsen
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to captionRefer to caption
(a) image produced by Dreambooth of Greg Rutkowski
(b) image produced by Dreambooth of Even Amundsen
(c) mix of Greg Rutkowski and Even Amundsen models

### v2

@@ 1 Introduction
1. The artist earns more compensation the closer the generation aligns with their style. 2. Artist receives extra compensation if the generated content is similar to content they have previously created. 1. Transferring the style to new content not inherent to the given artist. 2. Combining the styles of multiple artists.
@@ 4.1 Evaluation of a single style
Thus, we form a dataset as depicted in Figure 4. We aim to assess the extent to which the images from the second and third columns influence the first one. As previously mentioned, we evaluate this influence by measuring the similarity score between the style-less generations, the original artworks, and the final result. This evaluation produces the graph shown in Figure 5(a). From this graph, it’s evident that the original artworks by the artist bear a significant resemblance to the final generation. This observation can also be visually corroborated using Figure 3. Now, it is imperative to ascertain the correct methodology for employing the available data to compute the Shapley Values. The foundational equation is delineated as follows:
$\phi_{i}(v)=\sum_{S\subseteq N\setminus\{i\}}\frac{|S|!\cdot(|N|-|S|-1)!}{|N|!}\left(v(S\cup\{i\})-v(S)\right)$ (1)
where $\phi_{i}(v)$ represents the Shapley value for player $i$, $N$ is the set of all players, $S$ is a subset of $N$ without $i$, $v(S)$ is the value function for coalition $S$, $|S|$ is the number of players in $S$, $|N|$ is the total number of players, and then we perform the sum over all $S\subseteq N\setminus\{i\}$.
We then adapt this equation to accommodate our available data. Each player is represented as one of two categories: a group of original artworks (training data) or a group of generated images (generated data). We consider the maximum score achievable by such a collaboration as a constant, set to 1. A collaboration that attains a score of 1 is the group of images resulting from a combination of all available data and styles. Other coalitions receive lower scores, which are evaluated based on the embedding similarity between those coalitions and the target coalition, which represents the most rich and diverse amalgamation of available data and models.
For instance, when considering the combination of two entities — ’Alena Aenami,’ representing an artist who supplies training data, and ’StableDiffusion-v1-5,’ functioning as a generative model — we observe the emergence of three distinct coalitions. Each coalition is associated with specific prompts that activate its respective contribution (in cases where activation is required):
’Alena Aenami’ functions as the provider of training data. ’SD’ acts as the source of generated content, with the prompt: ’a woman.’ ’SD + Alena Aenami’ represents a synergy of the generative model with Alena Aenami’s training data, utilizing the prompt: ’a woman, by Alena Aenami’. (see Figure 5(b)) Subsequently, we gather a selection of images from each collaborative effort, apply Equation  1 for analysis, and achieve the following outcomes: {’Alena Aenami’:0.7422, ’Stable Diffusion’:0.2578}. This suggests that in the event of a sale of an image generated by this model using the specified prompt, the proceeds should be apportioned between the developers of the generative model and the artist in such a ratio. This division is proposed to ensure that the collaborative effort is equitably recognized and that all contributing parties benefit commensurately.
@@
(a) prompt: a mech
(b) prompt: a mech, by Alena Aenami
@@ 4.2.1 Dreambooth
(c) image produced by Dreambooth of Greg Rutkowski
(d) image produced by Dreambooth of Leon Tukker
(e) mix of Greg Rutkowski and Leon Tukker models
(f) outputs from the SD1.5 model without any specific style applied.
@@
(c) image produced by Dreambooth of Greg Rutkowski
(d) image produced by Dreambooth of Even Amundsen
(e) mix of Greg Rutkowski and Even Amundsen models